OpenStudy (anonymous):

sqrt(2x+4)=sqrt(6x+6)

OpenStudy (cggurumanjunath):

(2x+4)=(6x+6) 2x-6x=6-4 -4x=2 x=-0.5

OpenStudy (anonymous):

Thanks!

OpenStudy (cggurumanjunath):

wc